Suppose a perpetuity bond pays an interest of $40. The financial investor can earn a 5% return elsewhere. What is the maximum amount buyers would be willing to pay for this bond?

Note: I think we have to find the perpetual bond first. Then calculate the yield of the bond then compare to the promised rate of return of other investment assets. However, I am still confused on how to carry out this question and what the actual answer is
Answer #1

This question asks us to find the principal for which both bond yield and the alternative investment option equate, meaning:

p*0.05=40

or p=40*20=800

Thus, 800 is the maximum amount that a buyer shall be willing to pay for this bond